# Tony Lindeberg

### v5

@@ 6.1 Extensions of the approach
[…]
@@
Such an approach is in a straightforward way compatible with learning of the scale parameters by back propagation, based on automatic differentiation in deep learning environments. It would be conceptually very straightforward to extend the theoretical framework and the experimental evaluations presented in this paper to incorporating also a detailed analysis of these two additional classes of discrete derivative approximations for Gaussian derivative operators of hybrid type. For reasons of space constraints, we have, however, not been able to include corresponding in-depth analyses of those additional discretization methods here. [^13] In particular, regarding the theoretical properties of these hybrid discretization approaches, it should be mentioned that, due to the approximation of the spatial derivative operators $\partial_{x^{\alpha}}$ by the central difference operators $\delta_{x^{\alpha}}$, in combination with the unit $l_{1}$-norm normalization of the corresponding spatial smoothing operations, the hybrid derivative approximation kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}$ and $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (116) and (117) will obey similar response properties (77) and (78) to monomial input, as the previously treated approach, based on the combination of spatial smoothing using the discrete analogue of the Gaussian kernel with central difference operators $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64). In other words, the hybrid kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ and $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (116) and (117) will with $p_{k}(x)=x^{k}$ obey $T_{\scriptsize\mbox{hybr-sampl},x^{M}}(\cdot;\;s)*p_{M}(\cdot)=M!$ and $T_{\scriptsize\mbox{hybr-int},x^{M}}(\cdot;\;s)*p_{M}(\cdot)=M!$, as well as $T_{\scriptsize\mbox{hybr-sampl},x^{M}}(\cdot;\;s)*p_{N}(\cdot)=0$ and $T_{\scriptsize\mbox{hybr-int},x^{M}}(\cdot;\;s)*p_{N}(\cdot)=0$ for $M>N$. In these respects, these hybrid discretization approaches $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ and $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ could be expected to constitute better approximations of Gaussian derivative operators at very fine scales, than their corresponding non-hybrid counterparts, $T_{\scriptsize\mbox{normsampl},x^{\alpha}}(n;\;s)$ and $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\;s)$ according to (19) and (20). Simultaneously, these hybrid approaches will also be computationally much more efficient than their non-hybrid counterparts, in situations where Gaussian derivatives of multiple orders are to be computed at the same scale. For very small values of the scale parameter, the spatial smoothing with the normalized sampled Gaussian kernel $T_{\scriptsize\mbox{normsampl},x^{\alpha}}(n;\;s)$ can, however, again be expected to lead to systematically too small amounts of spatial smoothing (see Figure 3). For larger values of the scale parameter, the spatial smoothing with the integrated Gaussian kernel $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\;s)$ would, however, be expected to lead to a scale offset (see Figure 3), influenced by the variance of a box filter over each pixel support region. Furthermore, these hybrid approaches will not be guaranteed to obey information reducing properties from finer to coarser levels of scale, in terms of either non-creation of new local extrema, or non-enhancement of local extrema, as the discrete analogues of Gaussian derivatives $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64) obey. In situations, where the modified Bessel functions of integer order are immediately available, the approach, based on the combination of spatial smoothing with the discrete analogue of the Gaussian kernel with central differences, should therefore be preferable in relation to these hybrid approaches. In situations, where the modified Bessel functions of integer order are, however, not available as full-fledged primitive in a deep learning framework, these hybrid approaches could, on the other hand, be considered as interesting alternatives to the regular sampled or integrated Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\;s)$ and $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\;s)$, because of their substantially better computational efficiency, in situations when spatial derivatives of multiple orders are to be computed at the same scale. What remains to explore, is how these hybrid discretization approaches compare to the previously treated three main classes of discretization methods, with respect to the set of quantitative performance measures defined and then evaluated experimentally in Sections 3.73.8 and Sections 4.44.5, as well as with respect to integration in different types of computer vision and/or deep learning algorithms.

### v7

@@ 6.1 Extensions of the approach
Concerning extensions of the approach, with regard to applications in deep learning, for which the modified Bessel functions $I_{n}(s)$, underlying the definition of the discrete analogue of the Gaussian kernel $T_{\scriptsize\mbox{disc}}(n;\;s)$ according to (26), are currently generally not available in standard frameworks for deep learning, a possible alternative approach consists of instead replacing the previously treated sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\;s)$ according to (53) or the integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\;s)$ according to (54) by the families of hybrid discretization approaches obtained by: (i) first smoothing the image with either the normalized sampled Gaussian kernel $T_{\scriptsize\mbox{normsampl}}(n;\;s)$ according to (19) or the integrated Gaussian kernel $T_{\scriptsize\mbox{int}}(n;\;s)$ according to (20), and then applying central difference operators $\delta_{x^{\alpha}}$ of the form (60) to the spatially smoothed data.
@@
Such an approach is in a straightforward way compatible with learning of the scale parameters by back propagation, based on automatic differentiation in deep learning environments. It would be conceptually very straightforward to extend the theoretical framework and the experimental evaluations presented in this paper to incorporating also a detailed analysis of these two additional classes of discrete derivative approximations for Gaussian derivative operators of hybrid type. For reasons of space constraints, we have, however, not been able to include corresponding in-depth analyses of those additional discretization methods here. [^13] In particular, regarding the theoretical properties of these hybrid discretization approaches, it should be mentioned that, due to the approximation of the spatial derivative operators $\partial_{x^{\alpha}}$ by the central difference operators $\delta_{x^{\alpha}}$, in combination with the unit $l_{1}$-norm normalization of the corresponding spatial smoothing operations, the hybrid derivative approximation kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}$ and $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (116) and (117) will obey similar response properties (77) and (78) to monomial input, as the previously treated approach, based on the combination of spatial smoothing using the discrete analogue of the Gaussian kernel with central difference operators $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64). In other words, the hybrid kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ and $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (116) and (117) will with $p_{k}(x)=x^{k}$ obey $T_{\scriptsize\mbox{hybr-sampl},x^{M}}(\cdot;\;s)*p_{M}(\cdot)=M!$ and $T_{\scriptsize\mbox{hybr-int},x^{M}}(\cdot;\;s)*p_{M}(\cdot)=M!$, as well as $T_{\scriptsize\mbox{hybr-sampl},x^{M}}(\cdot;\;s)*p_{N}(\cdot)=0$ and $T_{\scriptsize\mbox{hybr-int},x^{M}}(\cdot;\;s)*p_{N}(\cdot)=0$ for $M>N$. In these respects, these hybrid discretization approaches $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ and $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ could be expected to constitute better approximations of Gaussian derivative operators at very fine scales, than their corresponding non-hybrid counterparts, $T_{\scriptsize\mbox{normsampl},x^{\alpha}}(n;\;s)$ and $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\;s)$ according to (19) and (20). Simultaneously, these hybrid approaches will also be computationally much more efficient than their non-hybrid counterparts, in situations where Gaussian derivatives of multiple orders are to be computed at the same scale. For very small values of the scale parameter, the spatial smoothing with the normalized sampled Gaussian kernel $T_{\scriptsize\mbox{normsampl},x^{\alpha}}(n;\;s)$ can, however, again be expected to lead to systematically too small amounts of spatial smoothing (see Figure 3). For larger values of the scale parameter, the spatial smoothing with the integrated Gaussian kernel $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\;s)$ would, however, be expected to lead to a scale offset (see Figure 3), influenced by the variance of a box filter over each pixel support region. Furthermore, these hybrid approaches will not be guaranteed to obey information reducing properties from finer to coarser levels of scale, in terms of either non-creation of new local extrema, or non-enhancement of local extrema, as the discrete analogues of Gaussian derivatives $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64) obey. In situations, where the modified Bessel functions of integer order are immediately available, the approach, based on the combination of spatial smoothing with the discrete analogue of the Gaussian kernel with central differences, should therefore be preferable in relation to these hybrid approaches. In situations, where the modified Bessel functions of integer order are, however, not available as full-fledged primitive in a deep learning framework, these hybrid approaches could, on the other hand, be considered as interesting alternatives to the regular sampled or integrated Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\;s)$ and $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\;s)$, because of their substantially better computational efficiency, in situations when spatial derivatives of multiple orders are to be computed at the same scale. What remains to explore, is how these hybrid discretization approaches compare to the previously treated three main classes of discretization methods, with respect to the set of quantitative performance measures defined and then evaluated experimentally in Sections 3.73.8 and Sections 4.44.5 (see (Lindeberg 2024) for a treatment about this topic) as well as with respect to integration in different types of computer vision and/or deep learning algorithms.